# Tur[k]ingBench: A Challenge Benchmark for Web Agents

### v2

@@ 3.6 Evaluation Protocol
Our evaluation protocol consistent of a loop over the various choices of evaluation tasks. Specifically means that, the evaluation script provide the model with the URL where the task can be reached. Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
Furthermore, in this paper we also provide the list of field names on which we will evaluate model responses. This arguable simplifies the task a bit, however, our benchmarks leaves it open for any future analysis with open-ended interaction without any additional hand-holding about the names of the target fields.
@@
To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 121 tasks that can be used for supervised model development. (ii) Test: a set of 20 tasks that be used for evaluating model quality. (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3. In this version of the work, we do not provide any experiments on Test since they require more sophisticated engineering that is beyond the scope of this work and should be addressed in future work.
@@ 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
We benchmark various models with architectural choices on TurkingBench. We note that, our goal here is to provide reasonable baselines for our proposed benchmark and hence, we avoid specialized models that may use specific assumptions or supervision of our data.
@@
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art models like GPT-4 and Claude-2.1.666Accessed via their APIs in January 2024. For GPT-4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few lines of HTML above and under the input field the model is currently solving. To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation task. This adds up to web pages evaluated with a total of roughly input fields evaluated.
We note that, besides these general-purpose models there are more specialized text-based models that are specialized for web-exploration either through pre-training on text data (Aghajanyan et al., 2021; 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2022; Chen et al., 2022) or visual perception of web (Dosovitskiy et al., 2020; Rust et al., 2022; Lee et al., 2022). We delegate such explorations that are likely to yield better results to the future.
An oracle model for ceiling performance.
We implement an oracle baseline that mimics the gold labels for each input. This data provides granular action sequences that facilitate future model developments. These action sequences are extracted by connecting crowdworker annotations to the name of the corresponding HTML input elements in task web pages and executing an appropriate action similar to the ones shown in Figure 5. A pseudo-code of the oracle model is shown in Alg.2.777See the a screencast of the oracle model’s execution. We note that, the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) that were included in Test (discussed in our evaluation split §3.5).
@@
We also evaluate a trivial baseline that executes nothing (hence the name, “do-nothing”). As we see later in the results, there are tasks that score more that zero. This is because there are tasks that ask, for example, corrections on a given text where at times doing nothing is the right action. Despite such settings, we will see the overall there is a big margin between the do-nothing baseline and the oracle baseline.
@@ 4.1 Empirical Results
We outline our empirical observations on our proposed benchmark.
Modalities |
Score (%) | |||||||
| Do-nothing | – | – | – | – | 7.8 | |||||||
| GPT4 | T | relevant | 7 | 5k | 21.3 | |||||||
| T | full | 7 | 82k | 39.3 | ||||||||
| GPT4-V | T+ V | relevant | 7 | 8.4k | 24.5 | |||||||
| T+ V | full | 7 | 86k | 41.7 | ||||||||
| Claude2.1 | T | full | 7 | 82k | 22.6 | |||||||
| Oracle | – | – | – | – | 100.0 |
@@
We show the results of evaluating our main models in Table 4. For each row, we indicate whether the input to the model contains text-only (T) vs. text-vision (T+V). All the models show have 7 in-context demonstrations of tasks and desired action. Note in the table we also show the size input prompts measured in subwords. The best performance is obtained by GPT-4 vision-language model (T+V) with access to full text of each task. This configuration also happens to have a extremely large prompt length (86k subwords) and it shows the remarkable ability of this model to exploit long-range dependencies. We note that the gains of the vision model (T+V) over text-only model (T) is minimal ( vs. ).
@@
For our top configurations in Table 4 (GPT-4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6. On 12 evaluation tasks (out of 20) the performance of these two models are notably different from one another. While it is intuitive to speculate that task with richer task interface likely benefit more from visual, in practice it was not clear what exact details details about task design leads to such differences.
@@
Performance across field type.
To better understand the difficulty of the task as a function of field types, we show the performance of GPT-4 (text input; 7 demonstrations) per input field type in Figure 8. We distinguish between “text” and “text area” because they’re defined with different tags in HTML (<input type="text"> vs. <textarea>). The former is often for short inputs and the latter for longer, multi-sentence outputs. This is also refelcted in lower performance results for “textarea” inputs.
5 Conclusion
We introduced TurkingBench to facilitate research on general-purpose web-based agents. Our benchmark is focused on tasks defined in the context of web pages, such as those that tend to appear on crowdsourcing platforms. Our benchmark includes a comprehensive Python-based framework that facilitates evaluation and model development. We hope this benchmark facilitates further development of web-based assistant agents.
@@ 5 Conclusion
There are various limitations in our work. (i) Despite the complexity of crowdsourcing tasks, they make up just a tiny fraction of the web. We are willing to make this trade-off to obtain natural tasks with ground-truth labels. Other existing benchmarks often need to develop artificial (manually-annotated) tasks and manually annotate gold labels for them. (ii) Because of the choice of domain, our tasks do not involve interactions that require jumping over multiple web pages—instead, all the tasks are contained within a given web page. Despite these simplifications, as we see in our experiments, the task remains challenging for the models. (iii) Our evaluation setup provides some hand-holding for the models by providing the names of the fields that need to be modified. Future work should study variants of our task where such hints are not provided to the model.
Author Contributions
KX led the major implementation of the evaluation framework and obtaining the final results. YK implemented the first version of the repository. KS, YW, AB participated in meetings, provided feedback and helped with the draft. JZ helped with the paper draft. DK lead the project, implementation and writing.
@@ Acknowledgements
This project is supported by a generous gift from Amazon. The authors would like to thank Chris Callison-Burch, Mark Dredze, Karthik Narasimhan, Nikhil Sharma, Owen Bianchi, and Elizabeth Salesky for their constructive discussions. GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).
6 Ethics Statement
A notable portion of humans spend more time on the web than the outside world (Perrin & Kumar, 2019). Consider an academic individual who spend long hours of their day responding to emails, writing proposals, scheduling talks, setting up meetings, etc. As a result, AI-driven automation of web navigation has the potential to simplify the experience of many and make web more accessible (Deng, 2023).
We acknowledge a concern that this work may enable technologies that would replace crowd workers, who have been an essential part of the AI development lifecycle. The concern is not specific to this work but applies broadly to all AI developments. We note that all the models discussed here are mainly enabled by developments outside this benchmark. Furthermore, the experimental results on our benchmark show that state-of-the-art self-supervised models are still far from the benchmark’s ceiling performance, even on our (simplified) evaluation setup. This means it will take much more progress before AI systems can fully solve autonomous crowdsourcing.
We hope this work enables benevolent use-cases of AI, such as AI systems that help improve the quality and productivity of crowd workers.

### v3

@@ 3.6 Evaluation Protocol
Our evaluation protocol consistent of a loop over the various choices of evaluation tasks. Specifically means that, the evaluation script provides the model with the URL where the task can be reached. Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
Based on this evaluation protocol, it should be evident that our benchmark does not require navigation between web pages, as noted in the limitations section. However, our benchmark does require multiple rounds of interaction to solve a given task. This is because nearly all of our tasks involve more than one step (input field), each needing to be addressed in a different round of interaction.
To reduce the difficulty of the task for our models, we provide the list of field names on which we will evaluate model responses. While this arguably simplifies the task, our benchmark remains open for future analysis with open-ended interaction, without any additional guidance on the names of the target fields.
@@
To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 125 tasks that can be used for supervised model development. (ii) Test: a set of 16 tasks that be used for evaluating model quality (names shown in the caption of Fig. 6). (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3. In this version of the work, we do not provide any experiments on Test since they require more sophisticated engineering that is beyond the scope of this work and should be addressed in future work.
@@ 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
We benchmark various models with different architectures on TurkingBench. Our goal is to provide reasonable baselines for our proposed benchmark, so we avoid specialized models that rely on specific assumptions or supervision of our data.
@@
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024. For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al. (2024). Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
We note that, besides these general-purpose models, there are more specialized text-based models for web exploration, either through pre-training on text data (Aghajanyan et al., 2021, 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2023; Chen et al., 2022), or visual perception of the web (Dosovitskiy et al., 2021; Rust et al., 2023; Lee et al., 2022; Kil et al., 2024). We leave such explorations, which are likely to yield better results, to future work.
Encoding the tasks for evaluation.
When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving. To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks. This adds up to web pages evaluated with a total of roughly input fields evaluated.
Modalities |
Score (%) | ||||||||||||
Do-nothing | – | – | – | – | – | – | 7.8 | ||||||||||||
open-source models | Llama3.1-Instruct | 8B | T | relevant | 3 | 1k | 128k | 23.2 | |||||||||||
8B | T | relevant | 7 | 5k | 128k | 25.0 | |||||||||||||
[…]
13B | T+ V | relevant | 3 | 1.6k | 4096 | 19.8 | |||||||||||||
[…]
76B | T+ V | relevant | 3 | 1.6k | 8192 | 30.4 | |||||||||||||
closed-source models | Claude2.1 | Unknown | T | full | 7 | 82k | 200k | 22.6 | |||||||||||
[…]
T | relevant | 3 | 4k | 128k | 18.4 | ||||||||||||||
T | relevant | 7 | 5k | 128k | 21.3 | ||||||||||||||
[…]
T | full | 3 | 28k | 128k | 35.7 | ||||||||||||||
T | full | 7 | 82k | 128k | 39.3 | ||||||||||||||
[…]
T+ V | relevant | 3 | 1.6k | 128k | 30.3 | ||||||||||||||
T+ V | relevant | 7 | 8k | 128k | 30.19 | ||||||||||||||
[…]
T+ V | full | 3 | 30k | 128k | 41.1 | ||||||||||||||
T+ V | full | 7 | 86k | 128k | 41.7 | ||||||||||||||
Oracle | – | – | – | – | – | – | 100.0 |
An oracle for ceiling performance.
We implement an oracle baseline that mimics the gold labels for each input. This oracle agent ensures the functional correctness of our evaluation. Our end-to-end evaluation process (model output execution lookup answers from web pages evaluation) is significantly more complex than a typical NLP benchmark, and any of these steps can fail. Achieving 100% accuracy with the oracle (ensuring functional correctness) took the lead student several months of effort. The oracle baseline essentially replicates the action sequences of crowdworkers for each HTML input element in the task web pages, executing appropriate actions similar to those shown in Figure 5. A pseudo-code of the oracle model is shown in Pseudocode 2.
We note that the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) included in Test (discussed in our evaluation split §3.5). A future use case of this oracle baseline can be to obtain granular action sequences for supervising models.
@@
We evaluate a trivial baseline that performs no action (no actions (hence, “do-nothing”). As we see in the results, this baseline scores more than zero because on some tasks doing nothing is the right action (e.g., making grammatical corrections to a given text that happens to be grammatical).
@@ 4.1 Empirical Results
We present the results of evaluating our main models in section 4. We experimented with models of varying parameter sizes. For each row, we indicate whether the model input contains text-only (T) or text-vision (T+V). Additionally, the table shows the size of input prompts measured in GPT-2 subwords Radford et al. (2019). Wherever possible, we evaluate the models with 7 in-context demonstrations of tasks and desired actions. The open-source vision-language models (T+V) had much shorter context windows, so we evaluated them using the “relevant” portion of the HTML code for the task or in-context demonstrations.
@@
The best performance is obtained by GPT4 vision-language model (T+V) with access to full text of each task. This is in-line with other recent observations Zheng et al. (2024a). This configuration also happens to have a extremely large prompt length (86k subwords) and it shows the remarkable ability of this model to exploit long-range dependencies. We note that the gains of the vision model (T+V) over text-only model (T) is minimal ( vs. ).
Open-source models rivaling GPT4.
Llama3.1-Instruct (8B params) notably outperforms GPT4 (text-only), achieving a score of 25.0% compared to GPT4’s 21.3% with 7 demonstrations when “relevant” bits of the input HTML are supplied. This result is particularly impressive given that Llama3.1-Instruct operates with significantly fewer parameters (8B) than GPT4’s rumored size. The comparison highlights Llama3.1-Instruct’s ability to efficiently leverage the provided context, potentially making it better suited for certain tasks despite its smaller size. Additionally, InternVL2 demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML. InternVL2’s score of 31.0% is remarkably close to GPT4’s 33%, despite the substantial difference in parameter sizes. This near-equivalence underscores the potential of InternVL2 to compete with proprietary models, offering similar performance with a more compact architecture.
Other open-source models, such as LLaVA-VL-Vicuna/Mistral (7B and 13B parameters), show varying degrees of success, with performance scores ranging from 13.8% to 22.7%. While these models are still behind Llama3.1 and InternVL2, their results demonstrate the diverse capabilities and potential of open-source approaches.
@@
For our top configurations in section 4 (GPT4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6. On 9 evaluation tasks (out of 16) the performance of these two models are notably different from one another. While it is intuitive to speculate that task with richer task interface likely benefit more from visual, empirical this it was not obvious what exact details details about task design leads to such differences.
@@
Performance across field types.
To better understand the task difficulty based on field types, we show the performance of GPT4 (text input; 7 demonstrations) for each input field type in Figure 8. We distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs. <textarea>). The former is typically used for short inputs, while the latter is for longer, multi-sentence outputs. A notable finding was the surprisingly lower performance of GPT4 compared to InternVL2 on “text” input fields, which we could not explain. However, since the data is skewed toward checkboxes (Figure 3), GPT4 performs better in the aggregate evaluations.
4.2 Error Analysis
We conducted human annotations to better understand the results in section 4. This analysis uses predictions from GPT4o (7 demos and “full” encoding) as it is one of the highest-performing models. One of the authors reviewed one instance from each of the 16 tasks, totaling 144 responses to the input fields.
Evaluating GPT4 responses.
Our annotator directly evaluated 134 (out of 144) responses that we successfully parsed by our evaluation metric (i.e., no parsing error). The annotator agreed with GPT4o’s responses in 60% (80 out of 134) of the cases, reaffirming that our benchmark remains challenging for the models.
Our analysis revealed several recurring issues. One common problem occurred in binary classification tasks, such as determining whether a word is an adjective with a similar meaning to a given word. For example, GPT4o incorrectly classified “discernible” as describing “modernity,” despite the lack of synonymy.
Another significant discrepancy was observed in tasks requiring GPT4o to generate both correct and incorrect answers. Instead of providing actual incorrect answers, GPT4o sometimes returned placeholders like //incorrect options//, failing to meet the task’s requirements. In some cases, GPT4o also made syntax errors. However, the evaluation focused on the content of the responses rather than their syntactical correctness, so syntax errors did not affect the assessment of answer accuracy. We would like to provide more details upon having more space.
5 Conclusion and Future Work
We introduced TurkingBench to facilitate research on web-based agents. Our benchmark focuses on tasks defined within the context of web pages, such as those commonly found on crowdsourcing platforms. It includes a comprehensive Python-based framework that supports both evaluation and model development. We hope this benchmark will drive further advancements in the development of web-based assistant agents.
Future work should explore modeling improvements for better web agents. For instance, a RAG-style approach could semantically chunk web pages into meaningful segments, a non-trivial task. Another avenue could involve using these agents as CoPilots for human annotation on crowdsourcing platforms, helping workers identify potential mistakes. We consider these extensions somewhat orthogonal to the primary focus of this work and hope future research will address them.
@@ 5 Conclusion and Future Work
We discuss several limitations: (i) No navigation between pages: Our benchmark does not require navigation between web pages. While inter-page navigation is important, it is not the only challenge. Effective understanding and manipulation of each page, which is our focus, remains a significant challenge for web agents. Our benchmark still requires navigation within each page, involving multiple rounds of interaction to solve tasks with multiple input fields. As shown in our experiments, this remains challenging for the models. We accept this trade-off to obtain more natural tasks compared to most existing benchmarks. (ii) Simplified evaluation: This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation. For simplicity, our evaluation setup provides some guidance by specifying the names of the fields to be modified. Future work should explore variants of our experiments where such hints are not provided to the model.
Ethical considerations.
We recognize concerns that this work could lead to technologies replacing crowd workers, who are vital to AI development. This concern is not unique to our work extends to all AI. Our results show that state-of-the-art models are still far from fully automating crowdsourcing tasks, even with simplified evaluation. Therefore, we hope our work enables benevolent use-cases of AI such as enhancing the quality and productivity of crowd workers rather than replacing them.
@@ Acknowledgements
This project is partly supported by ONR grant (N00014-24-1-2089) and a generous gift from Amazon. The authors would like to thank Chris Callison-Burch, Mark Dredze, Karthik Narasimhan, Nikhil Sharma, Owen Bianchi, and Elizabeth Salesky for their constructive discussions. GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).